# Intention Knowledge Graph Construction for User Intention Relation Modeling

### v1

@@ Abstract
Understanding user intentions is challenging for online platforms. Recent work on intention knowledge graphs addresses this but often lacks focus on connecting intentions, which is crucial for modeling user behavior and predicting future actions. This paper introduces a framework to automatically generate an intention knowledge graph, capturing connections between user intentions. Using the Amazon m2 dataset, we construct an intention graph with 351 million edges, demonstrating high plausibility and acceptance. Our model effectively predicts new session intentions and enhances product recommendations, outperforming previous state-of-the-art methods and showcasing the approach’s practical utility. [^1] We will release all data and code for this paper after this paper is published.
@@
Jiaxin Baithanks:   Equal contribution.1, Zhaobo Wang11footnotemark: 12, Junfei Cheng1, Dan Yu1, Zerui Huang1, Weiqi Wang1, Xin Liu3, Chen Luo3, Qi He3, Yanming Zhu2, Bo Li1, and Yangqiu Song1 1CSE, Hong Kong University of Science and Technology 2CSE, Shanghai Jiaotong University 3 Amazon.com {jbai, jchengao, dyuao, zhuangeu, wwangbw, bli, yqsong}@cse.ust.hk, {w19990112, yzhu }@sjtu.edu.cn , {xliucr, cheluo, qih}@amazon.com
@@ 1 Introduction
Understanding user intentions behind behaviors poses a significant challenge for online platforms. Recent research has introduced the concept of intention knowledge graphs Yu et al. (2023, 2024) to address this issue. These graphs connect various user behaviors, such as co-purchasing Yu et al. (2023) and queries Yu et al. (2024), to their corresponding intentions expressed in natural language. intention knowledge graphs have proven valuable in applications like product recommendation Yu et al. (2023) and search relevance Yu et al. (2024).
Users often have specific intentions, such as preparing for a Halloween party. They may want to dress up as a vampire or werewolf and create a spooky atmosphere with decorations. However, these intentions are often only implicitly suggested by actions, such as the items shown in Figure 1 (A). Previous attempts at constructing intention knowledge graphs have focused on using large language models to generate explanations for behaviors Yu et al. (2023), utilizing sources like session history and query keywords Yu et al. (2024). For instance, analyzing the browsing history of items can reveal the intention to dress up for Halloween.
However, previous methods have not explored how intentions interrelate. Explicitly modeling these connections aligns with the shift towards System II  [^2] http://www.iro.umontreal.ca/~bengioy/AAAI-9feb2020.pdf reasoning, emphasizing logical and sequential reasoning. Research has shown the importance of this reasoning in online behaviors Kleinberg et al. (2022), where users focus on long-term rewards rather than unconscious browsing. Therefore, we aim to model explicit intention relations.
Our goal extends beyond understanding initial intentions to predict subsequent ones. For example, intending to buy a desk might lead to buying an office chair. This inference can improve user behavior modeling and recommendations. However, building these connections is unexplored.
@@
Commonsense knowledge is crucial for modeling intention relationships. For example, planning a Halloween party requires understanding that costumes and decorations are likely intentions. We propose using commonsense relations to describe temporal and causal intention connections. After identifying intentions from user sessions, we use a classifier to determine inferential connections, enhancing our understanding of intention relationships. For instance, Figure 1 (B) demonstrates the plausible co-occurrence relationship between dressing up and decorating intentions.
Using inferential commonsense relations can be challenging to generalize to unseen situations. Concepts like conceptualization and instantiation help generalize commonsense reasoning Wang et al. (2023b, a, c, 2024). We incorporate this into our knowledge graph, using models to conceptualize intentions into broader concepts, connecting related intentions. For example, conceptualizing the intention to create a vampire costume can involve concepts like costumes and Halloween, linking it to similar intentions.
@@
We introduce the Intention Generation, Conceptualization, and Relation Classification (IGC-RC) framework for constructing a commonsense knowledge graph from user behaviors to address the challenges. The framework involves three general steps: (1) Intention Generation: Using large language models, we generate intentions from user sessions. (2) Conceptualization: We conceptualize the generated intentions using conceptualization models. (3) Relation Classification: We generate commonsense statements from intention relations and verify their plausibility.
@@
We apply the IGC-RC framework to the Amazon M2 session-based recommendation dataset, constructing a 351-million-edge knowledge graph. This graph captures intention-level relations and shows high acceptance in human evaluations, demonstrating the framework’s effectiveness. Our constructed knowledge graph allows accurate prediction of intentions in new user sessions and improves session-based recommendation models, outperforming previous approaches. This underscores the practical utility of our framework.
In summary, our contributions are: (1) We model connections between users’ deliberate and logical mental processes using an intention knowledge graph; (2) We develop the Intention Generation, Conceptualization, and Relation Classification (IGC-RC) framework, integrating user data and large language models; (3) We construct a large, high-quality Relational Intention Graph (RIG) from the Amazon M2 dataset, achieving superior performance in session recommendations.
@@ 2 Related Work
2.1 Knowledge Graphs for E-Commerce
Knowledge graphs (KGs) are crucial for enhancing recommendation systems on e-commerce platforms due to their structured and extensible nature. The Amazon Product Graph (Zalmout et al., 2021) exemplifies this by aligning Amazon’s catalog with external KGs like Freebase (Bollacker et al., 2008), extracting numerous attributes across product types. This interconnected web of product data enables Amazon to offer personalized shopping experiences. Similarly, Alibaba’s developments, such as AliCG (Zhang et al., 2021), AliCoCo (Luo et al., 2020), and AliMeKG (Li et al., 2020), leverage extensive e-commerce data for improved product recommendations and search capabilities. These KGs focus on item properties and categories without directly addressing the user motivations and intentions behind decisions. FolkScope Yu et al. (2023) and COSMO Yu et al. (2024) introduce intention knowledge graphs, incorporating user intentions from large language models and query-buy relations. However, they lack direct relations between intentions, such as causality and temporality. Our approach includes these commonsense relations to model intention dynamics more effectively.
Commonsense knowledge graphs (CSKGs) capture the world’s commonsense knowledge and enable machines to reason like humans. Notable CSKGs like ConceptNet (Speer et al., 2017), Atomic (Sap et al., 2019; Hwang et al., 2021), Discos (Fang et al., 2021), and WebChild (Tandon et al., 2017) have structured commonsense knowledge but are limited in describing concrete user intentions and motivations, essential for understanding behavior on e-commerce platforms.
@@ 2.1 Knowledge Graphs for E-Commerce
2.2 Session Understanding
Recent methods have explored using session history to reflect user intentions and enhance recommendation systems. Sequence modeling approaches, including Recurrent Neural Networks (RNNs) and Convolutional Neural Networks (CNNs), have been employed to model session data (Hidasi et al., 2015; Li et al., 2017; Liu et al., 2018; Tang and Wang, 2018). Advancements in session-based recommendation systems have focused on Graph Neural Networks (GNNs) to capture better session transitions (Li et al., 2021; Guo et al., 2022; Huang et al., 2022). Wu et al. (2019a) first utilized GNNs to capture complex transitions using graph structures. Subsequent research incorporated position, target information, global context, and highway networks to enhance performance (Pan et al., 2020; Xia et al., 2021). However, existing methods often simplify user intentions by focusing on item features or statistical metrics, resulting in an inadequate understanding of user preferences. Our approach aims to address this by explicitly modeling user intentions within sessions.
@@ 3.1 Intention Generation
We utilize the Amazon M2 dataset Jin et al. (2023), focusing on the English subset with 1.2 million sessions, to generate user intentions. The session history, comprised of sequences of user-browsed products, is processed to extract attributes such as titles, descriptions, models, sizes, and colors. These details are compiled into a JSON file, serving as the input for our intention generation module.
Using GPT-3.5, we generate concise, informative, and diverse user intentions, totaling 4.3 million from the dataset. Our approach surpasses FolkScope by using an advanced language model and eliminating ConceptNet relation constraints, fostering a more diverse output. See Figure 3 in the appendix for our prompt designs.
Precedence Succession Simultaneous Cause Result
Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1
Vera 0.73 0.52 0.61 0.78 0.66 0.72 0.75 0.83 0.79 0.75 0.73 0.74 0.76 0.92 0.83
+ Fine-tuning 0.86 0.97 0.91 0.88 0.92 0.90 0.87 0.93 0.90 0.84 0.91 0.88 0.81 0.99 0.89
@@ 3.2 Intention Relation Classification
A significant challenge is linking user intentions meaningfully. These intentions, often event-based, rely on commonsense knowledge to establish connections, such as understanding that purchasing a gift usually precedes celebrating Christmas (Speer et al., 2017; Sap et al., 2019; Hwang et al., 2021; Fang et al., 2021; Tandon et al., 2017).
As shown in Table 11, we first convert candidate edges into assertions using a template-based approach to achieve this. Then, the plausibility estimation model Vera (Liu et al., 2023a) evaluates these assertions. Finally, we set a plausibility threshold to add high-plausibility edges to the knowledge graph. We conducted binary classification on relational edges. We conduct expert annotation to improve the performance of VERA-T5-Base model. Two experts independently annotated the data used to train the VERA model. Each intention pair received at least two independent annotations, with relationships classified according to the Penn Discourse Treebank 2.0 Annotation Manual Prasad et al. (2008). The annotated relationships included temporal sequence, causality, and synchronization. The initial inter-annotator agreement was 85%, and disagreements were resolved through discussion between the experts. Overall, they annotated 1000 positive examples and randomly sampled another 1000 negative samples to construct a binary classification dataset, keeping 1600 for training and 200 for validation and testing. Table 4 shows the performance of the fine-tuned vera model. We applied this model to randomly drawn intention pairs, retaining those with a Vera score over 0.9.
@@ 3.3 Intention Conceptualization
A conceptualization provides an abstract, simplified view of a selected part of the world, encompassing objects, concepts, and relationships Gruber (1993); Himanen et al. (2019). Intention conceptualization extracts abstract concepts from user intentions to represent them in a knowledge graph, aiding e-commerce recommendation systems by modeling user intention dynamics.
Recent studies Wang et al. (2023b, a, c, 2024) primarily address entity and event conceptualization, leaving user intention conceptualization as an open challenge. Two main approaches exist: crowd-sourced and large language model (LLM) annotations Wang et al. (2024). While crowd-sourcing offers high-quality results, it is costly and limited in coverage. Thus, we utilize a large language model for intention conceptualization.
To ensure diverse, unambiguous conceptualizations for machine integration, we design specific prompts to guide the LLM. Using the Meta-Llama-3-8B-Instruct model, we generate conceptualizations for each intention. The prompts used are detailed in Figure 4, and examples of the generated conceptualizations are shown in Table 3.
@@ 4.1 Human Evaluation
Table 5: The intention generation quality comparison with FolkScope.
@@
Annotators evaluate two aspects of the generated intentions: plausibility and typicality. Plausibility refers to the likelihood that an assertion is valid based on its properties, usages, and functions. Typicality measures how well an assertion reflects specific features influencing user behavior, such as informativeness and causality. For example, "they are used for Halloween parties" is more informative than "they are used for the same purpose." We collected annotations for 3,000 session-intention pairs, each evaluated by three annotators. The inter-annotator agreement scores were 0.91 for plausibility and 0.74 for typicality. We use the same annotation guidelines and criteria from the FolkScope paper, and this can ensure the same standard of annotations on plausibility and typicality. As shown in Table 5, our RIG surpasses FolkScope(Yu et al., 2023) in plausibility and typically. Moreover, as shown in Table 11, the model successfully captures particular, long-tail intentions such as "Relieve discomfort and soothe itching caused by hemorrhoids" and "Purchase unscented baby wipes for sensitive skin." These examples illustrate the model’s ability to understand and articulate context-specific user needs.
@@ 4.2 Automatic Evaluation
This study evaluates the task of linking unseen sessions to an existing intention graph. The dataset, comprising all session-intention edges, was split into training, validation, and test sets (8:1:1), with each session linked to 2-4 positive intentions. For evaluation, we applied random negative sampling to generate 30 candidate intentions per session, ranking the positive intentions among them. We compared large language models (Mistral-7B-Instruct-v0.1, Meta-Llama-3-8B-Instruct, and flan-t5-large) to our RIG model. Baselines use perplexity scores derived via proper prompting to rank intentions, while RIG employs an embedding-based approach. Specifically, we generate embeddings for sessions and intentions using a sentence model and use SASRec as a session encoder to compute session representations. As shown in Table 7, our RIG model achieves competitive accuracy (e.g., best Hit@10) while significantly outperforming LLMs in inference speed (3.01ms vs. 2,021–4,102ms), demonstrating its practicality for real-world applications requiring fast decision-making.
@@
To evaluate the performance of conceptualization prediction, we constructed a dataset of intention-concept pairs derived from our intention knowledge graph. This dataset was split into training, validation, and test sets using an 8:1:1 ratio. The test set consisted of 147,801 intention-concept pairs, which were used to perform ranking tasks with various methods. For the baseline models, we employed large language models (LLMs) to perform ranking in a generative manner. Specifically, we experimented with Mistral-7B-Instruct-v0.3, Meta-Llama-3-8B-Instruct, and flan-t5-xl. For each intention, a candidate pool containing both true and false concepts was constructed, and the concepts were ranked based on their generation order by the LLMs. Similar to the intention prediction task, we applied negative sampling for each intention to rank positive concepts among a pool of 500 candidates. Our proposed method relied on an embedding-based approach, leveraging a fine-tuned embedding model to generate embeddings for intentions and concepts. This fine-tuning ensured that intention embeddings were closer to their corresponding positive concept embeddings. As shown in Table 8, the results demonstrate that our conceptualization prediction method, based on the intention knowledge graph, achieves superior performance compared to LLMs regarding both ranking accuracy and inference time.
@@
Product Recovery Benchmark
We constructed a session-intention pair dataset in previous sessions, including product IDs, descriptions, and user intentions. Here, we construct a new benchmarking dataset of the product-intention pairs from all session-intention pairs by assuming the items within the session share the same intentions as the sessions. These product-intention pairs were randomly split into training, validation, and test sets with an 8:1:1 ratio. This dataset served as the basis for evaluating the ability of different methods to recover relevant intentions for given products. To ensure a fair comparison, we focused on 1,203 overlapping products between the test set of the product-intention pairs in RIG and Folkscope. For these overlapped products, we compared the intentions generated by Folkscope and RIG. Rankings and evaluations were conducted on this same set of overlapping products, enabling a direct and balanced comparison of intention quality between the two systems. This methodology ensured the evaluation results reflected each method’s capability to generate relevant and accurate intentions. Table 9 provides the detailed evaluation results. We used pre-computed intention and product embeddings as input to a Multi-Layer Perceptron (MLP) scoring model to train our model. The MLP was trained using Noise Contrastive Estimation (NCE) loss to distinguish between relevant and irrelevant intentions. We assessed the model’s performance on the test set during the evaluation phase by analyzing one positive sample against ten negative samples. Cosine similarity scores were computed for ranking intentions, enabling precise comparisons across methods. Finally, we compared our method with Folkscope under identical experimental settings. The results, presented in Table 9, demonstrate that our knowledge graph significantly outperforms Folkscope in recovering relevant intentions for products, highlighting the superior efficacy of our approach.

### v2

@@ Abstract
Understanding user intentions is challenging for online platforms. Recent work on intention knowledge graphs addresses this, but often lacks focus on connecting intentions, which is crucial for modeling user behavior and predicting future actions. This paper introduces a framework to automatically generate an intention knowledge graph, capturing connections between user intentions. Using the Amazon m2 dataset, we construct an intention graph with 351 million edges, demonstrating high plausibility and acceptance. Our model effectively predicts new session intentions and enhances product recommendations, outperforming previous state-of-the-art methods and showcasing the approach’s practical utility. [^1] We will release all data and code for this paper after this paper is accepted.
@@
Jiaxin Baithanks:   Equal contribution.1, Zhaobo Wang11footnotemark: 12, Junfei Cheng1, Dan Yu1, Zerui Huang1, Weiqi Wang1, Xin Liu3, Chen Luo3, Yanming Zhu2, Bo Li1, and Yangqiu Song1 1CSE, Hong Kong University of Science and Technology 2CSE, Shanghai Jiaotong University 3 Amazon.com {jbai, jchengao, dyuao, zhuangeu, wwangbw, bli, yqsong}@cse.ust.hk, {w19990112, yzhu }@sjtu.edu.cn , {xliucr, cheluo}@amazon.com
@@ 1 Introduction
Understanding user intentions is a fundamental challenge for online platforms seeking to optimize user experience. Recent advances in intention knowledge graphs have made significant strides in connecting user behaviors to their underlying intentions. These graphs link various behaviors, such as co-purchasing patterns Yu et al. (2023) and search queries Yu et al. (2024), to intentions expressed in natural language, proving valuable for applications ranging from product recommendation to search relevance optimization.
However, a critical limitation of existing approaches lies in their inability to model the relationships between different intentions. Consider a user preparing for Halloween—they may have multiple interconnected intentions: dressing as a vampire, creating spooky decorations, and hosting a party. While their browsing history (Figure 1 (A)) implicitly suggests these intentions, existing knowledge graphs fail to capture how these intentions relate to each other through temporal, causal, and conceptual links.
This gap aligns with Bengio’s distinction between System I and System II reasoning, where the latter emphasizes deliberate, logical, and sequential thinking. Research demonstrates that online shopping behaviors frequently reflect System II reasoning Kleinberg et al. (2022), with users making conscious decisions toward long-term goals rather than engaging in unconscious browsing. By explicitly modeling intention relationships, we can better capture the deliberate nature of user behavior and create more accurate predictive models.
Our research extends beyond merely understanding users’ initial intentions to predicting subsequent ones, for instance, inferring that someone who intends to buy a desk may soon need an office chair. This predictive capability has profound implications for user behavior modeling, session understanding, and recommendation systems. However, constructing these intention-to-intention relationships remains unexplored in current research.
@@
We identify two key mechanisms for modeling these relationships. First, commonsense knowledge is essential for understanding how intentions connect—planning a Halloween party naturally entails both costume selection and decoration preparation. We propose leveraging inferential commonsense relations to describe temporal sequences (before/after) and causal connections (because/as a result) between intentions. After identifying intentions from user sessions, we employ a classifier to determine these inferential connections, building a more comprehensive model of user intention dynamics. Figure 1 (B) illustrates this approach, showing the plausible co-occurrence relationship between costume and decoration intentions.
Second, we incorporate abstract conceptualization to improve generalization beyond observed intention pairs. Recent work demonstrates that conceptualization and instantiation enhance commonsense reasoning Wang et al. (2023b, a). We integrate this insight by using models to abstract intentions into broader concepts, establishing connections between semantically related intentions. For example, conceptualizing a "vampire costume" intention involves higher-level concepts like "costumes" and "Halloween," linking it to other costume-related intentions (Figure 1 (C)).
@@
To address these challenges, we introduce the Intention Generation, Conceptualization, and Relation Classification (IGC-RC) framework for constructing a comprehensive commonsense knowledge graph of user intentions. Our approach follows three key steps: (1) Intention Generation: We leverage large language models to generate plausible intentions from user session data, capturing the underlying goals driving observable behaviors. (2) Conceptualization: We abstract these intentions into higher-level concepts, facilitating connections between semantically related intentions. (3) Relation Classification: We generate and verify commonsense statements describing relationships between intentions, creating a structured graph of interconnected user goals.
@@
We apply the IGC-RC framework to the Amazon M2 session-based recommendation dataset, constructing a Relational Intention Graph (RIG) with 351 million edges. This graph captures rich intention-level relationships and demonstrates high plausibility in human evaluations. Our extensive experiments show that RIG enables accurate prediction of intentions in new user sessions and significantly enhances the performance of session-based recommendation models, outperforming previous state-of-the-art approaches.
Our contributions include:
  • We pioneer the modeling of connections between users’ deliberate mental processes through an intention knowledge graph that captures temporal, causal, and conceptual relationships between intentions.
  • We develop the IGC-RC framework, a novel methodology that integrates user behavior data with large language models to automatically construct rich, multi-faceted intention knowledge graphs.
  • We construct RIG, a large-scale, high-quality intention knowledge graph from the Amazon M2 dataset, demonstrating superior performance in session understanding and recommendation tasks.
By bridging the gap between discrete user behaviors and their interconnected intentions, our work represents a significant advancement in user modeling for e-commerce and recommendation systems, with broad implications for creating more responsive and intuitive online platforms.
@@ 2 Related Work
E-commerce platforms increasingly rely on knowledge graphs (KGs) for personalized experiences. Notable examples include Amazon Product Graph (Zalmout et al., 2021), which aligns with Freebase (Bollacker et al., 2008), and Alibaba’s ecosystems: AliCG (Zhang et al., 2021), AliCoCo (Luo et al., 2020), and AliMeKG (Li et al., 2020). While these KGs effectively represent item properties, they typically miss user intentions.
Recent work by FolkScope Yu et al. (2023) and COSMO Yu et al. (2024) incorporates user intentions through large language models, but lacks inter-intention relationships crucial for modeling goal-oriented behavior. Commonsense knowledge graphs like ConceptNet (Speer et al., 2017), ATOMIC (Sap et al., 2019; Hwang et al., 2021), Discos (Fang et al., 2021), and WebChild (Tandon et al., 2017) structure general knowledge but don’t address e-commerce-specific intentions.
Session-based recommendation has evolved from sequence modeling approaches using RNNs and CNNs (Hidasi et al., 2015; Li et al., 2017; Liu et al., 2018; Tang and Wang, 2018) to Graph Neural Networks (Li et al., 2021a; Guo et al., 2022; Huang et al., 2022). Wu et al. (2019a) pioneered GNNs for modeling session transitions, with subsequent improvements incorporating contextual information (Pan et al., 2020; Xia et al., 2021). However, these methods often simplify user intentions, whereas our approach models them explicitly within their relational context.
@@ 3 IGC-RC Framework
We introduce the Intention Generation, Conceptualization, and Relation Classification (IGC-RC) framework for constructing comprehensive intention knowledge graphs from user behavior data. Our approach integrates large language models with structured knowledge representation techniques to build a rich understanding of user intentions and their relationships.
@@ 3.1 Intention Generation
We leverage the Amazon M2 dataset Jin et al. (2023), specifically its English subset containing 1.2 million sessions. For each session, we extract product attributes including titles, descriptions, specifications, and technical details to create a comprehensive representation of user browsing history.
Using GPT-3.5, we generate 4.3 million diverse user intentions from these session histories. Our approach improves upon previous work by employing a more capable language model and removing restrictive relation constraints, resulting in more natural and contextually appropriate intentions. The prompting strategy (detailed in Appendix Figure 3) focuses on extracting concise verb phrases that capture user goals underlying their browsing behavior. Quantitative analysis in Appendix Figure 11 confirms significantly higher n-gram diversity compared to Folkscope because of using a better generation model and removal of conceptnet relation constraints.
Precedence Succession Simultaneous Cause Result Overall
Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1
Vera 0.73 0.52 0.61 0.78 0.66 0.72 0.75 0.83 0.79 0.75 0.73 0.74 0.76 0.92 0.83 0.75 0.73 0.74
+ Fine-tuning 0.86 0.97 0.91 0.88 0.92 0.90 0.87 0.93 0.90 0.84 0.91 0.88 0.81 0.99 0.89 0.85 0.94 0.89
@@ 3.2 Intention Relation Classification
Creating meaningful connections between intentions requires understanding their complex logical, temporal, and causal relationships. We develop a three-step approach to establish these connections: (1) Template-based assertion generation: We transform intention pairs into natural language assertions using templates that express five relationship types: precedence, succession, simultaneity, cause, and result (examples in Table 12). (2) Plausibility estimation: We employ the Vera model (Liu et al., 2023a), fine-tuned on expert-annotated data, to assess the plausibility of these assertions. The annotation process followed the Penn Discourse Treebank 2.0 guidelines Prasad et al. (2008), with two experts independently evaluating each sample (85% initial agreement). (3) Threshold-based edge selection: We retain only high-confidence relationships (Vera score > 0.9) to ensure graph quality. The fine-tuned model achieves 0.89 F1 score overall (Table 4), with particularly strong performance on temporal relationships.
This approach enables us to model diverse intention relationships including temporal sequence (before/after), co-occurrence (simultaneously), and causality (because/results in).
@@ 3.3 Intention Conceptualization
To improve generalization and abstraction, we develop techniques to map specific intentions to broader concepts. Building on theoretical foundations from knowledge representation Gruber (1993); Himanen et al. (2019) and recent advances in conceptualization Wang et al. (2023b, 2024), we create a novel approach specifically for intentions.
We employ Meta-Llama-3-8B-Instruct with carefully designed prompts (Appendix Figure 4) to generate concise, non-redundant concept sets for each intention. The prompt emphasizes three key qualities: (1) representativeness: concepts must accurately capture the intention’s essence; (2) unambiguity: concepts should have clear, focused meanings; and (3) complementarity: concepts should cover different semantic aspects of the intention.
Table 3 demonstrates the effectiveness of this approach, showing how diverse intentions are mapped to concise concept sets that capture their essential characteristics. These concepts serve as connective tissue in our knowledge graph, enabling us to link semantically related intentions that might otherwise appear unrelated based on surface features.
By integrating these three components, intention generation, relation classification, and conceptualization, our IGC-RC framework creates a rich, structured representation of user intentions and their relationships.
@@ 4.1 Human Evaluation
+ ConceptNet Rels 0.9542 0.6430
Table 5: The intention generation quality comparison with FolkScope. The comparison is only used to show the quality of intention generation, rather than the overall quality of two knowledge graphs.
@@
Annotators evaluate two aspects of the generated intentions: plausibility and typicality. Plausibility refers to the likelihood that an assertion is valid based on its properties, usages, and functions. Typicality measures how well an assertion reflects specific features influencing user behavior, such as informativeness and causality. For example, "they are used for Halloween parties" is more informative than "they are used for the same purpose." We collected annotations for 3,000 session-intention pairs, each evaluated by three annotators. The inter-annotator agreement scores were 0.91 for plausibility and 0.74 for typicality. We use the same annotation guidelines and criteria from the FolkScope paper, and this can ensure the same standard of annotations on plausibility and typicality. As shown in Table 5, our RIG surpasses FolkScope (Yu et al., 2023) in plausibility and typically. Moreover, as shown in Table 12, the model successfully captures particular, long-tail intentions such as "Relieve discomfort and soothe itching caused by hemorrhoids" and "Purchase unscented baby wipes for sensitive skin." These examples illustrate the model’s ability to understand and articulate context-specific user needs. Meanwhile, adding ConceptNet relations as constraints is not useful for plausibility also negatively impacts the intention typicality.
@@ 4.2 Automatic Evaluation
This study evaluates the task of linking unseen sessions to an existing intention graph. The dataset, comprising all session-intention edges, was split into training, validation, and test sets (8:1:1) based on sessions, with each session linked to 2-4 positive intentions. For evaluation, we applied random negative sampling to generate 30 candidate intentions per session, ranking the positive intentions among them. We compared large language models (Mistral-7B-Instruct-v0.1, Meta-Llama-3-8B-Instruct, and flan-t5-large) to our RIG model. Baselines use perplexity scores derived via proper prompting to rank intentions, while RIG employs an embedding-based approach. Specifically, we generate embeddings for sessions and intentions using a sentence model and use SASRec as a session encoder to compute session representations. As shown in Table 7, our RIG model achieves competitive accuracy (e.g., best Hit@10) while significantly outperforming LLMs in inference speed (3.01ms vs. 2,021–4,102ms), demonstrating its practicality for real-world applications requiring fast decision-making.
@@
To evaluate the performance of conceptualization prediction, we constructed a dataset of intention-concept pairs derived from our intention knowledge graph. This dataset was split into training, validation, and test sets using an 8:1:1 ratio. The test set consisted of 147,801 intention-concept pairs, which were used to perform ranking tasks with various methods. For the baseline models, we employed large language models (LLMs) to perform ranking in a generative manner. Specifically, we experimented with Mistral-7B-Instruct-v0.3, Meta-Llama-3-8B-Instruct, and flan-t5-xl. For each intention, a candidate pool containing both true and false concepts was constructed, and the concepts were ranked based on their generation order by the LLMs. Similar to the intention prediction task, we applied negative sampling for each intention to rank positive concepts among a pool of 500 candidates. Our proposed method relied on an embedding-based approach, leveraging a fine-tuned embedding model to generate embeddings for intentions and concepts. This fine-tuning ensured that intention embeddings were closer to their corresponding positive concept embeddings.
As shown in Table 8, the results demonstrate that our conceptualization prediction method, based on the intention knowledge graph, achieves superior performance compared to LLMs regarding both ranking accuracy and inference time.
@@
Product Recovery
We constructed a session-intention pair dataset in previous sessions, including product IDs, descriptions, and user intentions. Here, we construct a new benchmarking dataset of the product-intention pairs from all session-intention pairs by assuming the items within the session share the same intentions as the sessions. These product-intention pairs were randomly split into training, validation, and test sets with an 8:1:1 ratio. This dataset served as the basis for evaluating the ability of different methods to recover relevant intentions for given products. To ensure a fair comparison, we focused on 1,203 overlapping products between the test set of the product-intention pairs in RIG and Folkscope. For these overlapped products, we compared the intentions generated by Folkscope and RIG. Rankings and evaluations were conducted on this same set of overlapping products, enabling a direct and balanced comparison of intention quality between the two systems. This methodology ensured the evaluation results reflected each method’s capability to generate relevant and accurate intentions. Table 9 provides the detailed evaluation results. We used pre-computed intention and product embeddings as input to a Multi-Layer Perceptron (MLP) scoring model to train our model. The MLP was trained using Noise Contrastive Estimation (NCE) loss to distinguish between relevant and irrelevant intentions. We assessed the model’s performance on the test set during the evaluation phase by analyzing one positive sample against ten negative samples. Cosine similarity scores were computed for ranking intentions, enabling precise comparisons across methods. Finally, we compared our method with Folkscope under identical experimental settings. The results, presented in Table 9, demonstrate that our knowledge graph significantly outperforms Folkscope in recovering relevant products from intentions, highlighting the superior efficacy of our approach.